OpenStudy (anonymous):

How do you solve the rational equation 1/2 + 1/4 = 1/x?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (kaylaprincess):

1/2 + 1/4 = 3/4 = 1/(4/3) x = 4/3

OpenStudy (kaylaprincess):

( i answered your closed question, surry ) ^.^